OpenStudy (anonymous):

If you were to solve the following system by substitution, what would be the best variable to solve for and from what equation? 2x+6y=9 3x-12y=15 A. x, in the first equation B. x, in the second equation C.y, in the second equation D. y, in the first equation

OpenStudy (anonymous):

would anyone know how to solve this equation?

OpenStudy (anonymous):

I would look for an equation where you can make x or y = all whole values.

OpenStudy (anonymous):

meaning 3x-12y=15 would be x-4y=5??? after ....

OpenStudy (anonymous):

Look at the first equation 2x+6y=9..what happens if we try to determine the value for x...we have subtract 6y from both sides and we have 2x=9-6y..now we have to divide by 2 and we end up with x=9/2 - 3y...it works but because there is a fraction it is not the easiest method. This make sense?

OpenStudy (anonymous):

Now look at if we try to get a value for y in the first equation...what do you end up with...do the same with the second equation,

OpenStudy (anonymous):

mmmhhhhh okkkk then, thank you

OpenStudy (anonymous):

Welcome. Find the one equation that has no fractions,

OpenStudy (anonymous):

ok then n_n